# GraphBEV: Towards Robust BEV Feature Alignment for Multi-Modal 3D Object Detection

### v2

@@ 4 Experiments
In this section, we present the experimental setup of GraphBEV, and evaluate the performance of 3D object detection on nuScenes[4] dataset. Additionally, we followed the solution provided by Ref.[12] to simulate feature misalignment scenarios.
@@ 4.1 Experimental Setup
We evaluate GraphBEV on the challenging large-scale nuScenes dataset [4], which is collected with a 32-beam LiDAR and six cameras. The dataset is generally split into 700/150/150 scenes for training/validation/testing. The six images cover 360-degree surroundings and the dataset provides calibration matrices that enable precise projection from 3D points to 2D pixels. We use the mAP and NDS across all categories as the primary metrics for evaluation following [1, 34, 29]. Note that NDS metric is a weighted average of mAP and other breakdown metrics (e.g.,translation, scale, orientation, velocity, and attributes errors). In addition, to validate the robustness of feature alignment, we followed Ref. [12] to simulate misalignment that LiDAR and camera projection metric errors. It is worth noting that Ref. [12] only added noise to the validation dataset, without the train and test datasets.
@@
We implement GraphBEV within the PyTorch [39], built upon the open-source BEVFusion [34] and OpenPCDet [53]. For the LiDAR branch, feature encoding is performed using SECOND [61] to obtain LiDAR BEV features, with voxel dimensions set to [0.075m, 0.075m, 0.2m] and point cloud ranges specified as [-54m, -54m, -5m, 54m, 54m, 3m] across the X, Y, and Z axes, respectively. The camera branch employs a Swin Transformer [32] as the camera backbone, integrating Heads of numbers 3, 6, 12, 24, and utilizes FPN [16] for fusing multi-scale feature maps. The resolution of input images is adjusted and cropped to 256 $\times$ 704. In the LSS [40] configuration, frustum ranges are set with X coordinates [-54m, 54m, 0.3m], Y coordinates: [-54m, 54m, 0.3m], Z coordinates [-10m, 10m, 20m], and depth range: [1m, 60m, 0.5m].
During training, we employ data augmentation for 10 epochs, including random flips, rotations (within the range [$-\frac{\pi}{4}$, $\frac{\pi}{4}$]), translations (with std=0.5), and scaling in [0.9, 1.1] for LiDAR data enhancement. We use CBGS [74] to resample the training data. Additionally, we use random rotation in [$-5.4^{\circ}$, $5.4^{\circ}$] and random resizing in [0.38, 0.55] to augment the images. The Adam optimizer [20] is used with a one-cycle learning rate policy, setting the maximum learning rate to 0.001 and weight decay to 0.01. The batch size is 24, and training is conducted on 8 NVIDIA GeForce RTX 3090 24G GPUs. During inference, we remove Test Time Augmentation (TTA) data augmentation, and the batch size is set to 1 on an A100 GPU. All latency measurements are taken on the same workstation with an A100 GPU.
@@ 4.2 Comparisons with State-of-the-Art Methods
Table 1: Comparison with the SOTA methods on the nuScenes validation and test set. ‘C.V.’, ‘Motor.’, ‘Ped.’, and ‘T.C.’ are short for construction vehicle, motorcycle, pedestrian, and traffic cone, respectively. The Modality column: ‘L’ = only LiDAR data, ‘LC’ = the use of both LiDAR and camera data. ${}^{\dagger}$ means using TTA (test-time augmentation). The best performances are marked with bold font.
@@
Table 2: Comparison with the SOTA methods on BEV map segmentation on nuScenes validation set. The Modality column: ‘L’ = only LiDAR data, ‘LC’ = the use of both LiDAR and camera data.
@@
As shown in Table 1, our GraphBEV outperforms the baseline BEVFusion[34] by +1.6% in mAP and +1.5% in NDS. In general, the open-source nuScenes dataset has minimal feature alignment issues but cannot entirely avoid them, as illustrated in Figure 1(a). Specifically, our GraphBEV utilizes the KD-Tree algorithm to search the neighbor depth around the projected depth provided by LiDAR-to-camera. Compared to BEVFusion[34], our GraphBEV demonstrates significant improvements in small objects, such as Barrier by +4.0%, Bike by +2.2%, and Pedestrian by +2.7%. It is primarily caused by small objects being more sensitive to feature misalignment, where even slight misalignment between LiDAR and camera can result in greater misalignment in small objects. Furthermore, our GraphBEV and ObjectFusion tackle the issue of feature misalignment in BEV-based multi-modal methods [34, 29]. While ObjectFusion[5] introduces a post-fusion paradigm utilizing RoI Pooling fusion, our GraphBEV does not alter the paradigm of BEVFusion[34, 29], slightly outperforms ObjectFusion by 0.7% in mAP and 0.3% in NDS. Additionally, our GraphBEV achieved SOTA performance when we submitted the detection results for the nuScenes test set to the official evaluation server. It surpassed the baseline BEVFusion[34] by 1.5% in mAP and 0.7% in NDS. Overall, we were able to address feature misalignment issues in BEVFusion[34, 29] without altering its excellent paradigm.
@@
To evaluate 3D object detection, we also assessed the generalization capability in the BEV Map Segmentation (semantic segmentation) task on the nuScenes [4] validation set, as shown in Table 2. Following the same training strategy as the baseline BEVFusion [34], we conducted evaluations within the [-50m, 50m]×[-50m, 50m] region around the ego car for each frame. We reported Intersection-over-Union (IoU) scores for drivable area, pedestrian cross, walkway, stop line, car park, and divider. Significant improvements were observed for drivable area, pedestrian cross, walkway, stop line, and car park, with only a minor decrease observed for divider. Overall, our GraphBEV demonstrates not only significant performance in 3D object detection but also strong generalization capability in BEV Map Segmentation.
@@ 4.3 Ablation Study
Table 3: Roles of Different Modules in GraphBEV for Feature Alignment on nuScenes validation set under clean setting and noisy misalignment setting. ‘C.V.’, ‘Motor.’, ‘Ped.’, and ‘T.C.’ are short for construction vehicle, motorcycle, pedestrian, and traffic cone, respectively. ‘+L (only)’ indicates the addition of only the LocalAlign module, ‘+G (only)’ indicates the addition of only the GlobalAlign module. GraphBEV denotes the addition of both LocalAlign and GlobalAlign modules. ‘LT(ms)’ represents latency. All latency measurements are conducted on the same workstation with an A100 GPU.
@@
Table 4: Robustness to weather conditions, different ego distances, different sizes on nuScenes [4] clean validation set. It is important to note that the evaluation metric is mAP (%).
Method Different Weather Conditions Different Ego Distances Different Object Sizes
Sunny Rainy Day Night Near Middle Far Small Moderate Large
Baseline [34] 68.2 69.9 68.5 42.8 77.5 60.9 34.8 44.7 54.5 60.4
GraphBEV 70.1 70.2 69.7 45.1 78.6 65.3 42.1 55.4 58.3 63.1
+1.9 +0.3 +1.2 +2.3 +1.1 +4.4 +7.3 +10.7 +3.8 +2.7
@@
To analyze the impact of misalignment, we conducted comparative experiments between our GraphBEV and BEVFusion [34]. It is noteworthy that in Table 3, we introduced misalignment in the nuScenes validation set, rather than in the training and testing sets, as done in Ref. [12]. We trained on the clean nuScenes [4] train dataset and evaluated performance under both clean and noisy misalignment settings. In the clean setting, GraphBEV outperforms BEVFusion [34] significantly, while in the noisy setting, the performance improvement is substantial. Furthermore, it is evident that BEVFusion [34] exhibits a significant decrease in metrics such as mAP and NDS from clean to noisy, whereas GraphBEV demonstrates a smaller decrease in performance metrics.
Notably, adding the LocalAlign or GlobalAlign module to BEVFusion [34] has minimal impact on latency compared to BEVFusion [34] , and the latency is lower than that of TransFusion[1] . When only the LocalAlign module is added to BEVFusion [34] , using the KD-Tree algorithm to build proximity relationships and fusing projected depth with neighbor depth to prevent feature misalignment, there are significant enhancements observed in both the clean and noisy misalignment settings. Adding only the GlobalAlign module to BEVFusion [34] also leads to noticeable improvements. Particularly, simultaneous addition of the LocalAlign and GlobalAlign modules shows strong performance in both clean and noisy settings.
@@
4.3.2 Robustness to Weather Conditions, Ego Distances and Object Sizes.
As shown in Table 4, we present a robustness analysis of our GraphBEV concerning different weather conditions. Various weather conditions influence the 3D object detection task. Following the approach of BEVFusion[34] and ObjectFusion [5], we partition the scenes in the validation set into the sunny, rainy, day, and night conditions. We outperform BEVFusion[34] under different weather conditions, especially in night scenes. Overall, our GraphBEV improves performance in sunny weather by accurate feature alignment and enhances performance in bad weather conditions.
Furthermore, we analyzed the impact of different ego distances and object sizes on the performance of GraphBEV. We categorized annotation and prediction ego distances into three groups: Near (0-20m), Middle (20-30m), and Far (>30m), and summarize the size distributions for each category, defining three equal-proportion size levels: Small, Moderate, and Large. It is evident that GraphBEV demonstrates significant performance improvements for distant and small objects. Compared to BEVFusion[34], our GraphBEV consistently enhances performance across all ego distances and object sizes, further narrowing performance gaps. Overall, our GraphBEV exhibits greater robustness to changes in ego distances and object sizes.
@@
As shown in Table 5, to analyze the impact of the hyperparameter $K_{\text{graph}}$ in the LocalAlign module on feature misalignment, we studied its effects under noisy misalignment settings on the nuScenes validation set. $K_{\text{graph}}$, as the number of nearest depths for LiDAR-to-camera projected depth in the LocalAlign module, influences the expressive capability of neighbor depth features. It is observed that our GraphBEV achieves optimal overall performance when $K_{\text{graph}}$ is 8. Therefore, selecting an appropriate $K_{\text{graph}}$ is crucial and may vary in other datasets. Furthermore, despite significant fluctuations in mAP due to changes in $K_{\text{graph}}$, overall performance surpasses that of BEVFusion.
@@ 5 Conclusion
In this work, we propose a robust fusion framework, GraphBEV, to address the feature misalignment issue for BEV-based methods [34, 29]. To mitigate feature misalignment caused by inaccuracy projected depth provided by LiDAR, we propose the LocalAlign module, which utilizes neighbor-aware depth features through Graph matching. Furthermore, to prevent global misalignment during the fusion of LiDAR and camera BEV features, we designed the GlobalAlign module to simulate offset noise, followed by aligning global multi-modal features through learnable offsets. Our GraphBEV significantly outperforms BEVFusion on the nuScenes validation set, especially in the noisy misalignment setting. It demonstrates that our GraphBEV greatly advances the application of BEVFusion in real-world scenarios, particularly in noisy misalignment.

### v3

@@ 4 Experiments
In this section, we present the experimental setup of GraphBEV and evaluate the performance of 3D object detection on the nuScenes[4] dataset. Additionally, we followed the solution provided by Ref.[12] to simulate feature misalignment scenarios.
@@ 4.1 Experimental Setup
We evaluate GraphBEV on the challenging large-scale nuScenes dataset [4], which was collected using a 32-beam LiDAR and six cameras. The dataset is typically split into 700/150/150 scenes for training, validation, and testing. The six images cover a 360-degree view of the surroundings, and the dataset provides calibration matrices that enable precise projection from 3D points to 2D pixels. We use mAP and NDS across all categories as the primary metrics for evaluation, following [1, 34, 29]. Note that the NDS metric is a weighted average of mAP and other breakdown metrics (e.g., translation, scale, orientation, velocity, and attribute errors). In addition, to validate the robustness of feature alignment, we followed Ref. [12] to simulate misalignment that LiDAR and camera projection metric errors. It is worth noting that Ref. [12] only added noise to the validation dataset, without the train and test datasets.
@@
We implement GraphBEV within the PyTorch [39], built upon the open-source BEVFusion [34] and OpenPCDet [54]. For the LiDAR branch, feature encoding is performed using SECOND [62] to obtain LiDAR BEV features, with voxel dimensions set to [0.075m, 0.075m, 0.2m] and point cloud ranges specified as [-54m, -54m, -5m, 54m, 54m, 3m] across the X, Y, and Z axes, respectively. The camera branch employs a Swin Transformer [32] as the backbone, integrating Heads of numbers 3, 6, 12, 24, and utilizing FPN [16] to fuse multi-scale feature maps. The resolution of input images is adjusted and cropped to 256 $\times$ 704. In the LSS [40] configuration, frustum ranges are set with X coordinates [-54m, 54m, 0.3m], Y coordinates: [-54m, 54m, 0.3m], Z coordinates [-10m, 10m, 20m], and depth range: [1m, 60m, 0.5m].
During training, we employ data augmentation for ten epochs, including random flips, rotations (within the range [$-\frac{\pi}{4}$, $\frac{\pi}{4}$]), translations (with std=0.5), and scaling in [0.9, 1.1] for LiDAR data enhancement. We use CBGS [76] to resample the training data. Additionally, we use random rotation in [$-5.4^{\circ}$, $5.4^{\circ}$] and random resizing in [0.38, 0.55] to augment the images. The Adam optimizer [20] is used with a one-cycle learning rate policy, setting the maximum learning rate to 0.001 and weight decay to 0.01. The batch size is 24, and training is conducted on 8 NVIDIA GeForce RTX 3090 24G GPUs. During inference, we remove Test Time Augmentation (TTA) data augmentation, and the batch size is set to 1 on an A100 GPU. All latency measurements are taken on the same workstation with an A100 GPU.
@@ 4.2 Comparisons with State-of-the-Art Methods
Table 1: Comparison with the SOTA methods on the nuScenes validation and test set. ‘C.V.’, ‘Motor.’, ‘Ped.’ and ‘T.C.’ are short for construction vehicles, motorcycles, pedestrians, and traffic cones. The Modality column: ‘L’ = only LiDAR data, ‘L.C.’ = using both LiDAR and camera data. means using TTA (test-time augmentation). The best performances are marked with bold font.
@@
Table 2: Comparison with the SOTA methods on BEV map segmentation on nuScenes validation set. The Modality column: ‘L’ = only LiDAR data, ‘LC’ = using both LiDAR and camera data.
@@
As shown in Table 1, our GraphBEV outperforms the baseline BEVFusion[34] by +1.6% in mAP and +1.5% in NDS. The open-source nuScenes dataset generally has minimal feature alignment issues but cannot entirely avoid them, as illustrated in Figure 1(a). Specifically, our GraphBEV utilizes the KD-Tree algorithm to search the neighbor depth around the projected depth provided by LiDAR-to-camera. Compared to BEVFusion[34], our GraphBEV demonstrates significant improvements in small objects, such as Barrier by +4.0%, Bike by +2.2%, and Pedestrian by +2.7%. It is primarily caused by small objects being more sensitive to feature misalignment, where even slight misalignment between LiDAR and the camera can result in greater misalignment in small objects. Furthermore, our GraphBEV and ObjectFusion tackle the issue of feature misalignment in BEV-based multi-modal methods [34, 29]. While ObjectFusion[5] introduces a post-fusion paradigm utilizing RoI Pooling fusion, our GraphBEV does not alter the paradigm of BEVFusion[34, 29], slightly outperforms ObjectFusion by 0.7% in mAP and 0.3% in NDS. Additionally, our GraphBEV achieved SOTA performance when we submitted the detection results for the nuScenes test set to the official evaluation server. It surpassed the baseline BEVFusion[34] by 1.5% in mAP and 0.7% in NDS. Overall, we could address feature misalignment issues in BEVFusion[34, 29] without altering its excellent paradigm.
@@
To evaluate 3D object detection, we also assessed the generalization capability in the BEV Map Segmentation (semantic segmentation) task on the nuScenes validation set, as shown in Table 2. Following the same training strategy as the baseline BEVFusion, we conducted evaluations within the [-50m, 50m]×[-50m, 50m] region around the ego car for each frame. We reported Intersection-over-Union (IoU) scores for drivable area, pedestrian crossing, walkway, stop line, car park, and divider. Significant improvements were observed for drivable area, pedestrian crossing, walkway, stop line, and car park, with only a minor decrease observed for the divider. Overall, our GraphBEV demonstrates not only significant performance in 3D object detection but also strong generalization capability in BEV Map Segmentation.
@@ 4.3 Ablation Study
Table 3: Roles of Different Modules in GraphBEV for Feature Alignment on nuScenes validation set under clean setting and noisy misalignment setting. ‘C.V.’, ‘Motor.’, ‘Ped.’ and ‘T.C.’ are short for construction vehicles, motorcycles, pedestrians, and traffic cones. ‘+L (only)’ indicates the addition of only the LocalAlign module, and ‘+G (only)’ indicates only the GlobalAlign module. GraphBEV denotes the addition of both LocalAlign and GlobalAlign modules. ‘L.T. (ms)’ represents latency. All latency measurements are conducted on the same workstation with an A100 GPU.
@@
To analyze the impact of misalignment, we conducted comparative experiments between our GraphBEV and BEVFusion [34]. It is noteworthy that in Table 3, we introduced misalignment into the nuScenes validation set, rather than in the training and testing sets, as follow Ref. [12]. We trained on the clean nuScenes [4] training dataset and evaluated performance under both clean and noisy misalignment conditions. In the clean setting, GraphBEV outperforms BEVFusion [34] significantly, while in the noisy setting, the performance improvement is substantial. Furthermore, it is evident that BEVFusion [34] exhibits a significant decrease in metrics such as mAP and NDS when transitioning from clean to noisy conditions, whereas GraphBEV demonstrates a smaller decrease in performance metrics.
Notably, adding the LocalAlign or GlobalAlign module to BEVFusion [34] has minimal impact on latency compared to BEVFusion [34] alone, and the latency is lower than that of TransFusion[1]. When only the LocalAlign module is added to BEVFusion [34] and the KD-Tree algorithm is used to build proximity relationships, while fusing projected depth with neighbor depth to prevent feature misalignment, significant enhancements are observed in both clean and noisy misalignment settings. Adding only the GlobalAlign module to BEVFusion [34] also leads to noticeable improvements. Particularly, the simultaneous addition of the LocalAlign and GlobalAlign modules exhibits strong performance in both clean and noisy settings.
@@
Table 5: Robustness to weather conditions on nuScenes [4] clean validation set. Notably, the evaluation metric is mAP.
Method Different Weather Conditions
Sunny Rainy Day Night
Baseline [34] 68.2 69.9 68.5 42.8
GraphBEV 70.1 70.2 69.7 45.1
Table 6: Robustness to different ego distances, different sizes on nuScenes [4] clean validation set. Notably, the evaluation metric is mAP.
Method Different Ego Distances Different Object Sizes
Near Middle Far Small Moderate Large
TransFusion-L [1] 77.5 60.9 34.8 44.7 54.5 60.4
Baseline[34] 79.4 64.9 40.0 50.3 58.7 64.0
GraphBEV 78.6 65.3 42.1 55.4 58.3 63.1
@@
As shown in Table 4, to analyze the impact of the hyperparameter $K_{\text{graph}}$ in the LocalAlign module on feature misalignment, we studied its effects under noisy misalignment settings on the nuScenes validation set. $K_{\text{graph}}$, which is the number of nearest depths for LiDAR-to-camera projected depth in the LocalAlign module, influences the expressive capability of neighboring depth features. It is observed that our GraphBEV achieves optimal overall performance when $K_{\text{graph}}$ is set to 8. Therefore, selecting an appropriate value for $K_{\text{graph}}$ is crucial and may vary across different datasets. Furthermore, despite significant fluctuations in mAP resulting from changes in $K_{\text{graph}}$, the overall performance still surpasses that of BEVFusion.
4.3.3 Robustness to Weather Conditions.
As shown in Table 5, we present a robustness analysis of our GraphBEV with respect to different weather conditions. Various weather conditions influence the 3D object detection task. Following the approach of BEVFusion[34] and ObjectFusion [5], we partition the scenes in the validation set into sunny, rainy, day, and night conditions. We outperform BEVFusion[34] under different weather conditions, especially in night scenes. Overall, our GraphBEV improves performance in sunny weather through accurate feature alignment and enhances performance in adverse weather conditions.
4.3.4 Robustness to Ego Distances and Object Sizes.
As shown in Table 6, we analyzed the impact of different ego distances and object sizes on the performance of GraphBEV. We categorized annotation and prediction ego distances into three groups: Near (0-20m), Middle (20-30m), and Far (>30m), and summarized the size distributions for each category, defining three equal-proportion size levels: Small, Moderate, and Large. It is evident that GraphBEV demonstrates significant performance improvements for distant and small objects. Compared to BEVFusion[34], our GraphBEV consistently enhances performance across all ego distances and object sizes, further narrowing the performance gaps. Overall, our GraphBEV exhibits greater robustness to changes in ego distances and object sizes.
@@ 5 Conclusion
In this work, we propose a robust fusion framework, GraphBEV, to address the feature misalignment issue for BEV-based methods. To mitigate feature misalignment caused by inaccurately projected depth provided by LiDAR, we propose the LocalAlign module, which utilizes neighbor-aware depth features through graph matching. Furthermore, to prevent global misalignment during the fusion of LiDAR and camera BEV features, we designed the GlobalAlign module to simulate offset noise, followed by aligning global multi-modal features through learnable offsets. Our GraphBEV significantly outperforms BEVFusion on the nuScenes validation set, especially in noisy misalignment settings. It demonstrates that our GraphBEV greatly advances the application of BEVFusion in real-world scenarios, particularly under noisy misalignment conditions.